If X is distributed normally with mean 12 and standard deviation 4. Find i) P(X>20) , ii) P(X<20), iii)

p(0<X<12)


1
Expert's answer
2020-08-06T17:12:52-0400

"\mu = 12"

"\sigma = 4"


Solution i) p(X>20)


"p(X>20) = 1 - p(X<20)"


"Z = \frac{x - \mu}{\sigma}" = "\frac{20-12}{4}" = 2


"p(X<20)" = probability at Z = 2


= 0.9772

"p(X > 20) = 1 - 0.9772"


Answer: 0.0228


Solution ii) p(X<20)


From (i) p(X<20) = 0.9772

= Probability at Z = 2


Answer: 0.9772


solution iii) p(0<X<12)

"p(0<X<12)" =

"p(x<12) - p(x<0)"


For p(X<12)


at


"Z = \frac{12-12}{4} = 0"


p(x<12) = 0.5


For p(x<0)


at


"Z=\frac{0-12}{4} = -3"

p(x<0) = 0.00135


"P(0<x<12) = 0.5 - 0.00135"


Answer: 0.49865